# Concentration Inequalities for Suprema of Empirical Processes with Dependent Data via Generic Chaining with Applications to Statistical Learning

### v1

@@ Abstract
This paper develops a general concentration inequality for the suprema of empirical processes with dependent data. The concentration inequality is obtained by combining generic chaining with a coupling-based strategy. Our framework accommodates high-dimensional and heavy-tailed (sub-Weibull) data. We demonstrate the usefulness of our result by deriving non-asymptotic predictive performance guarantees for empirical risk minimization in regression problems with dependent data. In particular, we establish an oracle inequality for a broad class of nonlinear regression models and, as a special case, a single-layer neural network model. Our results show that empirical risk minimzaton with dependent data attains a prediction accuracy comparable to that in the i.i.d. setting for a wide range of nonlinear regression models.
@@ 2 Basic Framework, Assumptions and Main Result
Let $\{\bm{Z}_{t}\}$ be a dependent sequence of random vectors where $\bm{Z}_{t}$ takes values in $\mathcal{Z}\subset\mathbb{R}^{d}$ for each $t$. Consider the class of functions $g(\bm{Z}_{t},\bm{\theta})$ indexed by $\bm{\theta}\in\Theta$. Our main objective consists in controlling the supremum of the empirical process associated with the average of the functions over a sequence of dependent random vectors of length $T$, that is
$\sup_{\bm{\theta}\in\Theta}\left|{1\over T}\sum_{t=1}^{T}g(\bm{Z}_{t},\bm{\theta})-\mathbb{E}(g(\bm{Z}_{t},\bm{\theta}))\right|~.$
@@
$\|g(\bm{Z}_{t},\bm{\theta}_{1})-g(\bm{Z}_{t},\bm{\theta}_{2})-\mathbb{E}(g(\bm{Z}_{t},\bm{\theta}_{1})-g(\bm{Z}_{t},\bm{\theta}_{2}))\|_{\psi_{\alpha}}\leq C_{\Theta}d_{\Theta}(\bm{\theta}_{1},\bm{\theta}_{2})~,$
and $(ii)$ for some ${\bm{\theta}}_{0}\in\Theta$ it holds that $\|g(\bm{Z}_{t},{\bm{\theta}}_{0})-\mathbb{E}g(\bm{Z}_{t},{\bm{\theta}}_{0})\|_{\psi_{\alpha}}\leq C_{\Theta}$.
@@
$\mathbb{E}\left(\sup_{\bm{\theta}\in\Theta}|g(\bm{Z}_{t},\bm{\theta})-g(\bm{Z}_{t}^{*},\bm{\theta})|\right)\leq\|d_{\mathcal{Z}}(\bm{Z}_{t},\bm{Z}_{t}^{*})\|_{L_{r}}~,$
@@
Talagrand’s functional is a measure of complexity of a class of functions (Talagrand, 2005). We say that a sequence of partition $\{\mathcal{A}_{k}\}_{k\geq 0}$ of $\Theta$ is admisible if the sequence is increasing [^1] An increasing sequence of partitions means that every set of $\mathcal{A}_{k+1}$ is included in a set of $\mathcal{A}_{k}$. and it is such that $|\mathcal{A}_{k}|\leq 2^{2^{k}}$ for $k=0,1,\ldots$. For any $\bm{\theta}\in\Theta$, denote by $A_{k}(\bm{\theta})$ the unique element of $\mathcal{A}_{k}$ that contains $\bm{\theta}$. Let $\Delta(A)$ denote the diameter of the set $A\subset\Theta$ associated with the distance $d_{\Theta}$. Finally, for $\alpha>0$ Talagrand’s functional $\gamma_{\alpha}$ is defined as
@@
$\sup_{\bm{\theta}\in\Theta}\left|{1\over T}\sum_{t=1}^{T}g(\bm{Z}_{t},\bm{\theta})-\mathbb{E}(g(\bm{Z}_{t},\bm{\theta}))\right|\geq C_{\alpha}C_{\Theta}\left({1+\gamma_{2}(\Theta)\over\sqrt{n}}\varepsilon_{1}^{1/2}+{1+\gamma_{\alpha}(\Theta)\over n^{1/\alpha\vee 1}}\varepsilon_{1}^{1/\alpha}\right)+C_{\mathcal{Z}}\varepsilon_{2}$
@@
$\sup_{\bm{\theta}\in\Theta}\left|{1\over T}\sum_{t=1}^{T}g(\bm{Z}_{t},\bm{\theta})-\mathbb{E}(g(\bm{Z}_{t},\bm{\theta}))\right|$
@@
holds at most with probability $13(T/n)\exp(-\varepsilon)$, where $C_{\alpha}$ is the same constant that appears in the statement of the theorem.
@@
Second, the supremum of the empirical process is bounded by three terms. The first two terms depend, respectively, on the Talagrand’s functionals $\gamma_{2}(\Theta)$ and $\gamma_{\alpha}(\Theta)$, which capture the complexity of the parameter space $\Theta$. When $\alpha\geq 2$ the first term dominates and we recover the classic sub-Gaussian concentration rate, with the effective sample size $n$ playing the role typically held by the (actual) sample size $T$ in the i.i.d. setting. On the contrary, when $\alpha<2$, the second term dominates, leading to a slower concentration rate, which is still controlled by the effective sample size $n$. The third term depends on the $\beta$-mixing coefficients and may be interpreted as a correction term arising from the fact that the sequence of random vectors is dependent rather than independent. It is important to highlight that the choice of effective sample size $n$ entails a trade-off (assuming that the $\beta$-mixing coefficients are decaying). The first two terms that depend on Talagrand’s functional are small when the effective sample size is large. On the contrary the third term that depends on the $\beta$-mixing coefficients is small when the effective sample size is also small.
@@
Fifth, it is important to highlight that the generic chaining proof requires some separability conditions on the empirical process to be satified (Boucheron et al., 2013, Ch. 11). In line with many authors, we assume throughout that these requirements are satisfied.
@@ 3 Application to Statistical Learning Theory
Consider the stationary time series $\{(Y_{t},\bm{X}_{t}^{\prime})^{\prime}\}$ where $(Y_{t},\bm{X}_{t}^{\prime})^{\prime}$ takes values in $\mathcal{Y}\times\mathcal{X}\subset\mathbb{R}\times\mathbb{R}^{d}$ which is assumed to be a closed set. We are interested in forecasting the prediction target $Y_{t}$ on the basis of the vector of predictors $\bm{X}_{t}$. The forecasts for the prediction target $Y_{t}$ are obtained from the class of prediction rules $f_{\bm{\theta}}:\mathcal{X}\rightarrow\mathcal{Y}$ indexed by $\bm{\theta}\in\Theta$. The square loss is used to measure prediction accuracy
$L(Y_{t},f_{\bm{\theta}}(\bm{X}_{t}))=(Y_{t}-f_{\bm{\theta}}(\bm{X}_{t}))^{2}~.$
@@
$\hat{\bm{\theta}}\in\arg\min_{\bm{\theta}\in\Theta}R_{T}(\bm{\theta})\text{ where }R_{T}(\bm{\theta})={1\over T}\sum_{t=1}^{T}(Y_{t}-f_{\bm{\theta}}(\bm{X}_{t}))^{2}~.$
@@
$R(\hat{\bm{\theta}})=\mathbb{E}((Y-f_{\hat{\bm{\theta}}}(\bm{X}))^{2}|\hat{\bm{\theta}}=\hat{\bm{\theta}}(\mathcal{D}))~,$ (3)
@@
A classic objective of statistical learning theory is to obtain a bound on the performance of the ERM relative to the optimal risk that can be achieved within the given class of prediction rules. Define $R(\bm{\theta})=\mathbb{E}\left((Y_{t}-f_{\bm{\theta}}(\bm{X}_{t}))^{2}\right)$. Our aim is to find a pair $(B_{T}(\Theta),\delta_{T})$ such that
@@
Let $\bm{Z}_{t}=(Y_{t},\bm{X}_{t}^{\prime})^{\prime}$ and define $g(\bm{Z}_{t},\bm{\theta})=(Y_{t}-f(\bm{X}_{t},\bm{\theta}))^{2}$. Then, we have
$\sup_{\bm{\theta}\in\Theta}\left|R_{T}(\bm{\theta})-R(\bm{\theta})\right|=\sup_{\bm{\theta}\in\Theta}\left|{1\over T}\sum_{t=1}^{T}g(\bm{Z}_{t},\bm{\theta})-\mathbb{E}(g(\bm{Z}_{t},\bm{\theta}))\right|~.$
@@
In order to apply Theorem 2.1 we assume that a number of high-level conditions hold. In Section 3.1 we verify that these conditions are satisified, for example, by a single-layer neural network model.
@@
and $(ii)$ $\|Y_{t}\|_{\psi_{2}}\leq C_{1}$ and $\|\sup_{\theta\in\Theta}f_{\bm{\theta}}(\bm{X}_{t})\|_{\psi_{2}}\leq C_{1}$,
@@
$R(\hat{\bm{\theta}})<\inf_{\bm{\theta}\in\theta}R(\bm{\theta})+C\left(\gamma_{2}(\Theta)\sqrt{\log(n)\over n}+\gamma_{1}(\Theta){\log(n)\over n}+C_{\mathcal{Z}}{1\over\sqrt{n}}\right)~,$
@@
holds at least with probability $1-13n^{-1}$. We recall that the rate of convergence $\sqrt{\log(n)/n}$ is typically referred to as the classical rate of convergence of empirical risk minimization in the learning literature with i.i.d. data (Devroye et al., 1996, Ch. 12). Thus, our results recovers the classical rate of converge once we replace the sample size $T$ with the effective sample size $n$. We highlight that the proposition relies on fairly weak conditions on the sequence of mixing coefficients. In particular, it requires (a sufficiently fast rate of) polynomial decay as opposed to several contributions in the literature which typically assume geometric decay (Jiang and Tanner, 2010; Brownlees and Gudmundsson, 2025; Brownlees and Llorens-Terrazas, 2025). It is worth noting that the faster the rate of decay of the mixing coefficients (as captured by a larger value of $\zeta$), the smaller the discrepancy between $n$ and $T$ (as reflected by a value of $\eta$ closer to unity). Finally, we note that the result stated in this proposition can be further sharpened. However, it is presented in its current form for illustrative purposes.
@@ 3.1 Single-layer Neural Network
In this section we detail the proof of Theorem 2.1. To simplify exposition throughout this section we use $g(\bm{Z}_{t},\bm{\theta})$ to denote $g(\bm{Z}_{t},\bm{\theta})-\mathbb{E}(g(\bm{Z}_{t},\bm{\theta}))$.
@@
$\sup_{\theta\in\Theta}\left|{1\over T}\sum_{t=1}^{T}g(\bm{Z}_{t},\bm{\theta})\right|=\sup_{\theta\in\Theta}{1\over T}\left|\sum_{t=1}^{T}g(\bm{Z}_{t},\bm{\theta})-\sum_{j=0}^{M-1}\sum_{i=0}^{n}g(\bm{W}^{*}_{i,j},\bm{\theta})-g(\bm{W}^{*}_{i,j},\bm{\theta})\right|$
$\quad=\sup_{\theta\in\Theta}{1\over T}\left|\sum_{j=0}^{M-1}\sum_{i=0}^{n}g(\bm{W}_{i,j},\bm{\theta})-\sum_{j=0}^{M-1}\sum_{i=0}^{n}g(\bm{W}^{*}_{i,j},\bm{\theta})+\sum_{j=0}^{M-1}\sum_{i=0}^{n}g(\bm{W}^{*}_{i,j},\bm{\theta})\right|$
$\quad\leq\sup_{\theta\in\Theta}{1\over Mn}\left|\sum_{j=0}^{M-1}\sum_{i=0}^{n}g(\bm{W}^{*}_{i,j},\bm{\theta})\right|+\sup_{\theta\in\Theta}{1\over Mn}\left|\sum_{j=0}^{M-1}\sum_{i=0}^{n}g(\bm{W}_{i,j},\bm{\theta})-\sum_{j=0}^{M-1}\sum_{i=0}^{n}g(\bm{W}^{*}_{i,j},\bm{\theta})\right|~.$
Note that the second equality follows from the fact that $g(\bm{W}_{i,j},\bm{\theta})=0$ when $\bm{W}_{i,j}=\bm{z}$. Furthermore, we have that
$\sup_{\theta\in\Theta}{1\over Mn}\left|\sum_{j=0}^{M-1}\sum_{i=0}^{n}g(\bm{W}^{*}_{i,j},\bm{\theta})\right|$
$\quad\leq{1\over Mn}\left|\sum_{j=0}^{M-1}\sum_{i=0}^{n}g(\bm{W}^{*}_{i,j},\bm{\theta}_{0})\right|+\sup_{\theta\in\Theta}{1\over Mn}\left|\sum_{j=0}^{M-1}\sum_{i=0}^{n}g(\bm{W}^{*}_{i,j},\bm{\theta})-g(\bm{W}^{*}_{i,j},\bm{\theta}_{0})\right|~,$
@@
$\mathbb{P}\left(\sup_{\bm{\theta}\in\Theta}\left|{1\over T}\sum_{t=1}^{T}g(\bm{Z}_{t},\bm{\theta})\right|\geq\varepsilon^{\prime}\right)\leq\mathbb{P}\left({1\over Mn}\sum_{j=0}^{M-1}\left|\sum_{i=0}^{n}g(\bm{W}^{*}_{i,j},\bm{\theta}_{0})\right|\geq{\varepsilon^{\prime}_{1}}\right)$
$\quad+\mathbb{P}\left(\sup_{\bm{\theta}\in\Theta}{1\over Mn}\sum_{j=0}^{M-1}\left|\sum_{i=0}^{n}g(\bm{W}^{*}_{i,j},\bm{\theta})-g(\bm{W}^{*}_{i,j},\bm{\theta}_{0})\right|\geq{\varepsilon^{\prime}_{2}}\right)$
$\quad+\mathbb{P}\left(\sup_{\bm{\theta}\in\Theta}{1\over Mn}\sum_{j=0}^{M-1}\left|\sum_{i=0}^{n}g(\bm{W}_{i,j},\bm{\theta})-g(\bm{W}^{*}_{i,j},\bm{\theta})\right|\geq{\varepsilon^{\prime}_{3}}\right)$
$\quad\leq{T\over n}\max_{0\leq j\leq M-1}\mathbb{P}\left(\left|{1\over n}\sum_{i=0}^{n}g(\bm{W}^{*}_{i,j},\bm{\theta}_{0})\right|\geq{\varepsilon^{\prime}_{1}}\right)$
$\quad+{T\over n}\max_{0\leq j\leq M-1}\mathbb{P}\left(\sup_{\bm{\theta}\in\Theta}\left|\sum_{i=0}^{n}g(\bm{W}^{*}_{i,j},\bm{\theta})-g(\bm{W}^{*}_{i,j},\bm{\theta}_{0})\right|\geq{\varepsilon^{\prime}_{2}}\right)$
$\quad+{T\over n}\max_{0\leq j\leq M-1}\mathbb{P}\left(\sup_{\bm{\theta}\in\Theta}\left|{1\over n}\sum_{i=0}^{n}g(\bm{W}_{i,j},\bm{\theta})-g(\bm{W}^{*}_{i,j},\bm{\theta})\right|\geq{\varepsilon^{\prime}_{3}}\right)~.$ (7)
@@
A.1theasm.(ii)(ii) and Proposition 4.2 imply that for any $j\in\{0,\ldots,M-1\}$ we have that
$\mathbb{P}\left(\left|{1\over n}\sum_{i=0}^{n}g(\bm{W}^{*}_{i,j},\bm{\theta}_{0})\right|\geq{C^{\prime}_{\alpha}C_{\Theta}\sqrt{(\varepsilon^{\prime}_{1})}\over\sqrt{n}}+{C^{\prime\prime}_{\alpha}C_{\Theta}(\varepsilon^{\prime}_{1})^{1\over\alpha}\over n^{1/\alpha\vee 1}}\right)\leq e\exp\left(-\varepsilon^{\prime}_{1}\right)~.$ (8)
Notice that in this result we are using the fact that the random variable $g(\bm{W}^{*}_{i,j},\bm{\theta})$ is degenerate at zero when $\bm{W}^{*}_{i,j}=\bm{z}$ and that in this case we have that $\|g(\bm{W}^{*}_{i,j},\bm{\theta})\|_{\psi_{\alpha}}<C_{\Theta}$.

### v2

@@ Abstract
This paper develops a general concentration inequality for the suprema of empirical processes with dependent data. The concentration inequality is obtained by combining generic chaining with a coupling-based strategy. Our framework accommodates high-dimensional and heavy-tailed (sub-Weibull) data. We demonstrate the usefulness of our result by deriving non-asymptotic predictive performance guarantees for empirical risk minimization in regression problems with dependent data. In particular, we establish an oracle inequality for a broad class of nonlinear regression models and, as a special case, a single-layer neural network model. Our results show that empirical risk minimizaton with dependent data attains a prediction accuracy comparable to that in the i.i.d. setting for a wide range of nonlinear regression models.
@@ 2 Basic Framework, Assumptions and Main Result
Let $\{g_{\bm{\theta}}:\bm{\theta}\in\Theta\}$ be a class of real-valued functions defined on $\mathcal{Z}\subset\mathbb{R}^{d}$ indexed by $\bm{\theta}\in\Theta$. Let $\{\bm{Z}_{t}\}$ be a dependent sequence of random vectors where $\bm{Z}_{t}$ takes values in $\mathcal{Z}$ for each $t$. Our main objective consists in controlling the supremum of the empirical process associated with the average of the functions $g_{\bm{\theta}}$ based on the a sequence $\{\bm{Z}_{1},\ldots,\bm{Z}_{T}\}$, that is
$\sup_{\bm{\theta}\in\Theta}\left|{1\over T}\sum_{t=1}^{T}g_{\bm{\theta}}(\bm{Z}_{t})-\mathbb{E}(g_{\bm{\theta}}(\bm{Z}_{t}))\right|~.$
@@
$\|g_{\bm{\theta}_{1}}(\bm{Z}_{t})-g_{\bm{\theta}_{2}}(\bm{Z}_{t})-\mathbb{E}(g_{\bm{\theta}_{1}}(\bm{Z}_{t})-g_{\bm{\theta}_{2}}(\bm{Z}_{t}))\|_{\psi_{\alpha}}\leq C_{\Theta}d_{\Theta}(\bm{\theta}_{1},\bm{\theta}_{2})~,$
and $(ii)$ for some ${\bm{\theta}}_{0}\in\Theta$ it holds that $\|g_{\bm{\theta}_{0}}(\bm{Z}_{t})-\mathbb{E}g_{\bm{\theta}_{0}}(\bm{Z}_{t})\|_{\psi_{\alpha}}\leq C_{\Theta}$.
@@
$\mathbb{E}\left(\sup_{\bm{\theta}\in\Theta}|g_{\bm{\theta}}(\bm{Z}_{t})-g_{\bm{\theta}}(\bm{Z}_{t}^{*})|\right)\leq\|d_{\mathcal{Z}}(\bm{Z}_{t},\bm{Z}_{t}^{*})\|_{L_{r}}~,$
@@
Talagrand’s functional is a measure of complexity of a class of functions (Talagrand, 2005). We say that a sequence of partition $\{\mathcal{A}_{k}\}_{k\geq 0}$ of $\Theta$ is admissible if the sequence is increasing [^1] An increasing sequence of partitions means that every set of $\mathcal{A}_{k+1}$ is included in a set of $\mathcal{A}_{k}$. and it is such that $|\mathcal{A}_{k}|\leq 2^{2^{k}}$ for $k=0,1,\ldots$. For any $\bm{\theta}\in\Theta$, denote by $A_{k}(\bm{\theta})$ the unique element of $\mathcal{A}_{k}$ that contains $\bm{\theta}$. Let $\Delta(A)$ denote the diameter of the set $A\subset\Theta$ associated with the distance $d_{\Theta}$. Finally, for $\alpha>0$ Talagrand’s functional $\gamma_{\alpha}$ is defined as
@@
$\sup_{\bm{\theta}\in\Theta}\left|{1\over T}\sum_{t=1}^{T}g_{\bm{\theta}}(\bm{Z}_{t})-\mathbb{E}(g_{\bm{\theta}}(\bm{Z}_{t}))\right|\geq C_{\alpha}C_{\Theta}\left({1+\gamma_{2}(\Theta)\over\sqrt{n}}\varepsilon_{1}^{1/2}+{1+\gamma_{\alpha}(\Theta)\over n^{1/\alpha\vee 1}}\varepsilon_{1}^{1/\alpha}\right)+C_{\mathcal{Z}}\varepsilon_{2}$
@@
$\sup_{\bm{\theta}\in\Theta}\left|{1\over T}\sum_{t=1}^{T}g_{\bm{\theta}}(\bm{Z}_{t})-\mathbb{E}(g_{\bm{\theta}}(\bm{Z}_{t}))\right|$
@@
holds at most with probability $13(T/n)\exp(-\varepsilon)$, where $C_{\alpha}$ is the same constant that appears in the statement of the theorem. This is obtained by choosing $\varepsilon_{1}=\varepsilon$ and $\varepsilon_{2}=\beta^{s/(r(r+s))}\left(\left\lfloor{T\over n+1}\right\rfloor\right)\exp(\varepsilon)$.
@@
Second, the supremum of the empirical process is bounded by three terms. The first two terms depend, respectively, on the Talagrand’s functionals $\gamma_{2}(\Theta)$ and $\gamma_{\alpha}(\Theta)$, which capture the complexity of the parameter space $\Theta$. When $\alpha\geq 2$ the first term dominates and we recover the classic sub-Gaussian concentration rate, with the effective sample size $n$ playing the role typically held by the (actual) sample size $T$ in the i.i.d. setting. On the contrary, when $\alpha<2$, the second term dominates, leading to a slower concentration rate, which is still controlled by the effective sample size $n$. The third term depends on the $\beta$-mixing coefficients and may be interpreted as a correction term arising from the fact that the sequence of random vectors is dependent rather than independent. It is important to highlight that the choice of effective sample size $n$ entails a trade-off (assuming that the $\beta$-mixing coefficients are decaying). The first two terms that depend on Talagrand’s functional are small when the effective sample size is large. On the contrary the third term that depends on the $\beta$-mixing coefficients is small when the effective sample size is also small.c
@@
Fifth, it is important to highlight that the generic chaining proof requires the empirical process to be separable, in the sense of the definition of (Boucheron et al., 2013, p. 305). In line with many authors, we assume throughout that these requirements are satisfied.
@@ 3 Application to Statistical Learning Theory
Consider the stationary time series $\{(Y_{t},\bm{X}_{t}^{\prime})^{\prime}\}$ where $Y_{t}$ takes values in $\mathcal{Y}\subset\mathbb{R}$ and $\bm{X}_{t}$ takes values in $\mathcal{X}\subset\mathbb{R}^{d}$, with both $\mathcal{Y}$ and $\mathcal{X}$ assumed to be closed sets. We are interested in forecasting the prediction target $Y_{t}$ on the basis of the vector of predictors $\bm{X}_{t}$. The forecasts for the prediction target $Y_{t}$ are obtained from the class of prediction rules $f_{\bm{\theta}}:\mathcal{X}\rightarrow\mathcal{Y}$ indexed by $\bm{\theta}\in\Theta$. The square loss is used to measure prediction accuracy
$L(Y_{t},f_{\bm{\theta}}(\bm{X}_{t})):=(Y_{t}-f_{\bm{\theta}}(\bm{X}_{t}))^{2}~.$
@@
$\hat{\bm{\theta}}\in\arg\min_{\bm{\theta}\in\Theta}R_{T}(\bm{\theta})\text{ where }R_{T}(\bm{\theta}):={1\over T}\sum_{t=1}^{T}(Y_{t}-f_{\bm{\theta}}(\bm{X}_{t}))^{2}~.$
@@
$R(\hat{\bm{\theta}}):=\mathbb{E}((Y-f_{\hat{\bm{\theta}}}(\bm{X}))^{2}|\hat{\bm{\theta}}=\hat{\bm{\theta}}(\mathcal{D}))~,$ (3)
@@
A classic objective of statistical learning theory is to obtain a bound on the performance of the ERM relative to the optimal risk that can be achieved within the given class of prediction rules. Define $R(\bm{\theta}):=\mathbb{E}\left((Y_{t}-f_{\bm{\theta}}(\bm{X}_{t}))^{2}\right)$. Our aim is to find a pair $(B_{T}(\Theta),\delta_{T})$ such that
@@
Let $\bm{Z}_{t}=(Y_{t},\bm{X}_{t}^{\prime})^{\prime}$ that takes values in $\mathcal{Z}=\mathcal{Y}\times\mathcal{X}$ and define $g_{\bm{\theta}}(\bm{Z}_{t})=(Y_{t}-f(\bm{X}_{t},\bm{\theta}))^{2}$. Then, we have
$\sup_{\bm{\theta}\in\Theta}\left|R_{T}(\bm{\theta})-R(\bm{\theta})\right|=\sup_{\bm{\theta}\in\Theta}\left|{1\over T}\sum_{t=1}^{T}g_{\bm{\theta}}(\bm{Z}_{t})-\mathbb{E}(g_{\bm{\theta}}(\bm{Z}_{t}))\right|~.$
@@
In order to apply Theorem 2.1, we assume that a number of high-level conditions hold. These conditions play the same role as A.1theasm and A.2theasm in the previous section, but are reformulated to fit the present framework. In Section 3.1 we verify that these conditions are satisfied, for example, by a single-layer neural network model.
@@
and $(ii)$ $\|Y_{t}\|_{\psi_{2}}\leq C_{1}$ and $\|\sup_{\theta\in\Theta}f_{\bm{\theta}}(\bm{X}_{t})\|_{\psi_{2}}\leq C_{1}$ .
@@
$R(\hat{\bm{\theta}})<\inf_{\bm{\theta}\in\theta}R(\bm{\theta})+C\left(\gamma_{2}(\Theta)\sqrt{\log(n)\over n}+\gamma_{1}(\Theta){\log(n)\over n}+{1\over\sqrt{n}}\right)~,$
@@
holds at least with probability $1-13n^{-1}$. We recall that the rate of convergence $\sqrt{\log(n)/n}$ is typically referred to as the classical rate of convergence of empirical risk minimization in the learning literature with i.i.d. data (Devroye et al., 1996, Ch. 12). Thus, our results recovers the classical rate of converge once we replace the sample size $T$ with the effective sample size $n$. We highlight that the proposition relies on fairly weak conditions on the sequence of mixing coefficients. In particular, it requires (a sufficiently fast rate of) polynomial decay as opposed to several contributions in the literature which typically assume geometric decay (Jiang and Tanner, 2010; Brownlees and Gudmundsson, 2025; Brownlees and Llorens-Terrazas, 2025). It is worth noting that the faster the rate of decay of the mixing coefficients (as captured by a larger value of $\zeta$), the smaller the discrepancy between $n$ and $T$ (as reflected by a value of $\eta$ closer to unity).
@@ 3.1 Single-layer Neural Network
In this section we detail the proof of Theorem 2.1. To simplify exposition throughout this section we use $g_{\bm{\theta}}(\bm{Z}_{t})$ to denote $g_{\bm{\theta}}(\bm{Z}_{t})-\mathbb{E}(g_{\bm{\theta}}(\bm{Z}_{t}))$.
@@
$\sup_{\theta\in\Theta}\left|{1\over T}\sum_{t=1}^{T}g_{\bm{\theta}}(\bm{Z}_{t})\right|=\sup_{\theta\in\Theta}{1\over T}\left|\sum_{t=1}^{T}g_{\bm{\theta}}(\bm{Z}_{t})-\sum_{j=0}^{M-1}\sum_{i=0}^{n}g_{\bm{\theta}}(\bm{W}^{*}_{i,j})-g_{\bm{\theta}}(\bm{W}^{*}_{i,j})\right|$
$\quad=\sup_{\theta\in\Theta}{1\over T}\left|\sum_{j=0}^{M-1}\sum_{i=0}^{n}g_{\bm{\theta}}(\bm{W}_{i,j})-\sum_{j=0}^{M-1}\sum_{i=0}^{n}g_{\bm{\theta}}(\bm{W}^{*}_{i,j})+\sum_{j=0}^{M-1}\sum_{i=0}^{n}g_{\bm{\theta}}(\bm{W}^{*}_{i,j})\right|$
$\quad\leq\sup_{\theta\in\Theta}{1\over Mn}\left|\sum_{j=0}^{M-1}\sum_{i=0}^{n}g_{\bm{\theta}}(\bm{W}^{*}_{i,j})\right|+\sup_{\theta\in\Theta}{1\over Mn}\left|\sum_{j=0}^{M-1}\sum_{i=0}^{n}g_{\bm{\theta}}(\bm{W}_{i,j})-\sum_{j=0}^{M-1}\sum_{i=0}^{n}g_{\bm{\theta}}(\bm{W}^{*}_{i,j})\right|~.$
Note that the second equality follows from the fact that $g_{\bm{\theta}}(\bm{W}_{i,j},\bm{\theta})=0$ when $\bm{W}_{i,j}=\bm{z}$. Furthermore, we have that
$\sup_{\theta\in\Theta}{1\over Mn}\left|\sum_{j=0}^{M-1}\sum_{i=0}^{n}g_{\bm{\theta}}(\bm{W}^{*}_{i,j})\right|$
$\quad\leq{1\over Mn}\left|\sum_{j=0}^{M-1}\sum_{i=0}^{n}g_{\bm{\theta}_{0}}(\bm{W}^{*}_{i,j})\right|+\sup_{\theta\in\Theta}{1\over Mn}\left|\sum_{j=0}^{M-1}\sum_{i=0}^{n}g_{\bm{\theta}}(\bm{W}^{*}_{i,j})-g_{\bm{\theta}_{0}}(\bm{W}^{*}_{i,j})\right|~,$
@@
$\mathbb{P}\left(\sup_{\bm{\theta}\in\Theta}\left|{1\over T}\sum_{t=1}^{T}g_{\bm{\theta}}(\bm{Z}_{t})\right|\geq\varepsilon^{\prime}\right)\leq\mathbb{P}\left({1\over Mn}\sum_{j=0}^{M-1}\left|\sum_{i=0}^{n}g_{\bm{\theta}_{0}}(\bm{W}^{*}_{i,j})\right|\geq{\varepsilon^{\prime}_{1}}\right)$
$\quad+\mathbb{P}\left(\sup_{\bm{\theta}\in\Theta}{1\over Mn}\sum_{j=0}^{M-1}\left|\sum_{i=0}^{n}g_{\bm{\theta}}(\bm{W}^{*}_{i,j})-g_{\bm{\theta}_{0}}(\bm{W}^{*}_{i,j})\right|\geq{\varepsilon^{\prime}_{2}}\right)$
$\quad+\mathbb{P}\left(\sup_{\bm{\theta}\in\Theta}{1\over Mn}\sum_{j=0}^{M-1}\left|\sum_{i=0}^{n}g_{\bm{\theta}}(\bm{W}_{i,j})-g_{\bm{\theta}}(\bm{W}^{*}_{i,j})\right|\geq{\varepsilon^{\prime}_{3}}\right)$
$\quad\leq{T\over n}\max_{0\leq j\leq M-1}\mathbb{P}\left(\left|{1\over n}\sum_{i=0}^{n}g_{\bm{\theta}_{0}}(\bm{W}^{*}_{i,j})\right|\geq{\varepsilon^{\prime}_{1}}\right)$
$\quad+{T\over n}\max_{0\leq j\leq M-1}\mathbb{P}\left(\sup_{\bm{\theta}\in\Theta}\left|\sum_{i=0}^{n}g_{\bm{\theta}}(\bm{W}^{*}_{i,j})-g_{\bm{\theta}_{0}}(\bm{W}^{*}_{i,j})\right|\geq{\varepsilon^{\prime}_{2}}\right)$
$\quad+{T\over n}\max_{0\leq j\leq M-1}\mathbb{P}\left(\sup_{\bm{\theta}\in\Theta}\left|{1\over n}\sum_{i=0}^{n}g_{\bm{\theta}}(\bm{W}_{i,j})-g_{\bm{\theta}}(\bm{W}^{*}_{i,j})\right|\geq{\varepsilon^{\prime}_{3}}\right)~.$ (7)
@@
We note that A.1theasm.(ii)(ii) and Proposition 4.2 imply that for any $j\in\{0,\ldots,M-1\}$ we have that
$\mathbb{P}\left(\left|{1\over n}\sum_{i=0}^{n}g_{\bm{\theta}_{0}}(\bm{W}^{*}_{i,j})\right|\geq{C^{\prime}_{\alpha}C_{\Theta}\sqrt{(\varepsilon^{\prime}_{1})}\over\sqrt{n}}+{C^{\prime\prime}_{\alpha}C_{\Theta}(\varepsilon^{\prime}_{1})^{1\over\alpha}\over n^{1/\alpha\vee 1}}\right)\leq e\exp\left(-\varepsilon^{\prime}_{1}\right)~.$ (8)
Notice that in this result we are using the fact that the random variable $g_{\bm{\theta}}(\bm{W}^{*}_{i,j})$ is degenerate at zero when $\bm{W}^{*}_{i,j}=\bm{z}$ and that in this case we have that $\|g_{\bm{\theta}}(\bm{W}^{*}_{i,j})\|_{\psi_{\alpha}}<C_{\Theta}$.